# The dimension of sparse random graph orders

### v1

@@ 6 Non-bipartite POS: proof of Theorem 11
Throughout this section, $p=c_{n}/n$ where $1\leq c_{n}\ll\log n$ and $q=1-p$. Set
@@
Proof of Theorem 11 By considering the dual of $P$ and Lemma 25, a.a.s. every component of the cover graph of $P[U[S]]$ is either a tree or unicyclic. By Theorems 3 and 4, a.a.s. $\dim P[D[S]],\dim P[U[S]]\leq 4$. Then, Theorem 8 implies that a.a.s. $\dim P\leq 2(8+\dim P[U\cup D]+\dim P[[n]\setminus S])$, where $U=U[S]\setminus S$ and $D=D[S]\setminus S$. By the definition of $S$, every element in $P[[n]\setminus S]$ is comparable to at most $10Kc_{n}\alpha=O(c_{n}^{2}e^{c_{n}})$ other elements. Thus, by Theorem 5, $\dim P[U\cup D]\leq\dim P[[n]\setminus S]=O(c_{n}^{4}e^{c_{n}})$. ∎
It only remains to prove Lemma 25. Unlike $P_{\mathcal{B}(n,n,p)}$ where the comparable graph and the cover graph are both $\mathcal{B}(n,n,p)$, the cover graph of $P_{\mathcal{G}(n,p)}$ is typically a proper subgraph of $\mathcal{G}(n,p)$, whereas the comparable graph of $P_{\mathcal{G}(n,p)}$ is a super graph of $\mathcal{G}(n,p)$ with a much greater density. To study $P_{\mathcal{G}(n,p)}$, it is much more convenient to consider the oriented version of $\mathcal{G}(n,p)$ where every edge $ij$ in $\mathcal{G}(n,p)$ is oriented from $i$ to $j$ if $i<j$. Since every graph on $[n]$ has a unique orientation we keep the same notation $\mathcal{G}(n,p)$ for its oriented version. Note that two elements $i$ and $j$ in $P_{\mathcal{G}(n,p)}$ are comparable if and only if there is a directed path joining $i$ and $j$ in $\mathcal{G}(n,p)$.
The key idea in the proof of upper-bounding $\dim P_{\mathcal{B}(n,n,p)}$ is that $S_{A}$ is much smaller than $A$, and thus, it is not likely to join vertices in $S_{A}$ by vertices in $B$ into a connected bicyclic subgraph. Situations in $P_{\mathcal{G}(n,p)}$ are quite different, as elements with large down-sets or up-sets form clusters. Suppose $v$ is an element with a large down-set. Then all the elements $u\in U_{P}(v)$ have a large down-set. Therefore, we expect many large cliques in the comparable graph of $P_{\mathcal{G}(n,p)}$, and they intersect in non-trivial ways. However, this does not mean that the cover graph of $P_{\mathcal{G}(n,p)}$ induced by the elements with large down-sets have complicated graph structures. Indeed, Lemma 25 shows that every component of the cover graph has at most two cycles. Unlike the study of $\dim P_{\mathcal{B}(n,n,p)}$, analysing the cover graph of $P[D[S]]$ for $P\sim P_{\mathcal{G}(n,p)}$ necessitates the identification and careful examination of suitable chosen structures.
@@
Proof of Observation 26 Suppose on the contrary that there exists an arc $(x,y)$ in $G_{D}[S]\setminus\mathcal{G}(V,p)$. It implies that there is a directed $x,y$-path $P_{xy}$ in $\mathcal{G}(n,p)$ such that $P_{xy}$ has length at least two, and none of the internal vertices is in $V$. On the other hand, there is a directed path $P_{yz}$ from $y$ to some $z\in S$ since $y$ is in the down-set of $S$. By considering the directed path obtained from concatenating $P_{xy}$ and $P_{yz}$, it follows that all the internal vertices on $P_{xy}$ must be in the down-set of $z$, and thus must be in $V$, leading to a contradiciton. ∎
It is easy to see that the cover graph of $P[D[S]\cup U[S]]$ does not have the property in Observation 26, which is one of the main reasons that we complete the proof of Theorem 11 by studying the dimension of $P[D[S]]$ and $P[U[S]]$, instead of studying $\dim P[D[S]\cup U[S]]$.
@@
To simplify the analysis, we hope to identify a set of directed paths as in Observation 27(b) that “behave somewhat independently”. We say the triple $(x,y,z)$ in $W(M)$ is a trilink if
@@
Choose $M$ such that the total number of maximal elements is minimized. Let ${\mathcal{T}}$ be a maximum collection of pairwise well-separated trilinks in $W(M)$, and for each trilink $(x,y,z)\in{\mathcal{T}}$, let $P_{y}$ be an arbitrary directed path from $y$ to some $s_{y}\in S$, whose existence is guaranteed by Observation 27(b). Let ${\mathcal{P}}$ be the collection of such directed paths. We demonstrate that our choice of $(M,{\mathcal{T}},{\mathcal{P}})$ satisfies (P1)–(P4).
(P1): it holds trivially when $t\leq 42$ by taking ${\mathcal{T}}=\emptyset$. Suppose that $t\geq 43$. Let $M^{\prime}$ be the vertex-induced subgraph of $W(M)$ obtained by deleting the red vertices in $W(M)$ and their neighbours. Since there were either two red vertices each of degree three, or a single red vertex with degree 4 in $W(M)$, it is easy to see that $M^{\prime}$ has at least $t-12$ edges. Moreover, $M^{\prime}$ is a collection of at most three disjoint paths (not necessarily directed paths), one of which — call it $P$ — must have length at least $(t-12)/3$. Since every edge in $M^{\prime}$ joins a vertex in $\overline{M}$ and a vertex in $\underline{M}$, there must exists at least $\lfloor((t-12)/3)/7\rfloor\geq t/21-2$ well-separated trilinks by taking a maximum collection of pairwise well-separated trilinks on $P$.
@@
Let $U\subseteq[n]$ and let $G_{\bar{U}}$ be the subgraph of $\mathcal{G}(n,p)$ induced by $[n]\setminus U$. Let $P_{\bar{U}}$ be the poset with cover graph $G_{\bar{U}}$.
@@
Proof of Lemma 29 For (a), obviously, $\mathbb{P}(|D_{P_{\bar{U}}}(u)|\geq k)\leq\mathbb{P}(|D_{P}(u)|\geq k)$, and we can obtain the bound in part (a) easily from (6). However, it follows even more straightforwardly from [5, Lemma 3.1 (iii)].
@@
    For every $u\in[n]$, $\mathbb{P}(|D_{P}[u]\cup U_{P}[u]|\geq\alpha),\mathbb{P}(|U_{P}[u]\cup U_{P}[u]|\geq\alpha)\leq(1+o(1))\exp(-Kc_{n})$.
@@
Proof of Lemma 25 Suppose $G_{D}[S]$ has a connected bicyclic subgraph $M$ and a collection ${\mathcal{P}}$ of dipaths satisfying all the properties in Lemma 28. By the claim below, we may assume that $W(M)$ has at least 100 vertices.
@@
We first bound the number of ways to construct $W(M)$, and then $M$. We define a family ${\mathcal{F}}$ of connected directed bicyclic graphs (who can possibly contain a double edge) that can be candidates for $W(M)$ as follows. The vertices are marked $+$, or --, or coloured red, so that they respond to the vertices in $\overline{M}$, $\underline{M}$, and the red vertices in $W(M)$.
@@
  • Each vertex can be marked $+$ or --, or unmarked;
@@
  • Every arc $(x,y)$ joins some $x$ marked -- and some $y$ marked $+$, unless $x$ or $y$ is red.
@@
Let $\Sigma_{1},\Sigma_{2}$ be disjoint sets. Let $R$ be a set (not necessarily disjoint from $\Sigma_{1}\cup\Sigma_{2}$) of cardinality between one and two, and let $a=|\Sigma_{1}\cup\Sigma_{2}\cup R|$. The number of graphs in ${\mathcal{F}}$ on $\Sigma_{1}\cup\Sigma_{2}\cup R$ such that all vertices in $\Sigma_{1}$ are marked $+$, all vertices in $\Sigma_{2}$ are marked --, and all vertices in $R$ are coloured red is at most $O(a^{4a})$.
@@
Explanation of (15).  The first sum is over all integers $h,\ell,a$ that denote the sizes of $\Sigma_{1}$, $\Sigma_{2}$ and the total number of vertices in $W(M)$. The second sum is over all possible choices of $(\Sigma_{1},\Sigma_{2},R)$ that is consistent with $(h,\ell,a)$, where $R$ denotes the set of red vertices in $W(M)$. The third sum is over all possible weights on the arcs of $W(M)$, which indicates the lengths of the directed paths in $M$ between two vertices of $W(M)$. The next sum is over all possible choices of $M$ that are consistent with $(\Sigma_{1},\Sigma_{2},R,{\boldsymbol{w}})$. The next sum is over all choices of $k$ well-separated trilinks. Given a set of ${{trilink}}$s ${\mathcal{T}}$, let ${\mathcal{N}}={\mathcal{N}}({\boldsymbol{w}})$ be the set of $j\in[k]$ such that $w^{(j)}<4Kc_{n}$ where $w^{(j)}$ is for $y_{j}$ where $(x_{i},y_{i},z_{i})\in{\mathcal{T}}$. Next, we sum over all choices of $s_{y_{j}}$ for $j\in{\mathcal{N}}({\boldsymbol{w}})$. The final sum is over all possible subsets $D_{j},U_{j},j\in{\mathcal{N}}$ that are pairwise disjoint, except that $D_{j}\cap U_{j}=y_{j}$ for all $j$. Recalling that $N({\boldsymbol{w}})=|{\mathcal{N}}({\boldsymbol{w}})|=|\{i:w^{(i)}<4Kc_{n}\}|$. By Claim 34(a), for each $j\in{\mathcal{N}}({\boldsymbol{w}})$ either $|D_{j}|\geq 4Kc_{n}\alpha$ or $|U_{j}|\geq 4Kc_{n}\alpha$. Thus, by Claim 34(b,c),
@@
×$\hskip 227.62204pt\times n^{N}2^{a}\Big{(}2n^{-1}\exp(-Kc_{n}/2)\Big{)}^{N}.$ (16)
Explanation of (16).  We call $j\in[a+1]$ heavy if $w_{j}\geq 2Kc_{n}$ and light otherwise. Since $N({\boldsymbol{w}})=N$, there are $k-N$ $y_{i}$s in a trilink in ${\mathcal{T}}$ such that $w^{(i)}\geq 4Kc_{n}$. Since each of these $y_{i}$s has degree two in $W(M)$, at least one of the directed path incident to $y_{i}$ in $M$ has length at least $2Kc_{n}$, i.e. the corresponding $j$ is heavy. Since each directed path joins two vertices in $M$, there must be at least $(k-N)/2$ heavy $j\in[a+1]$. Let ${\mathcal{A}}$ be the set of all possible assignments of “heavy” and “light” to $j\in[a+1]$ such that there are at least $(k-N)/2$ heavy elements. Obviously, $|{\mathcal{A}}|\leq 2^{a+1}$. Then,
@@
by choosing sufficiently large $K$ so that $28c_{n}e^{c_{n}-Kc_{n}/168}<1/$ 2. ∎
Proof of Claim 31 Fix $B,h,\ell$. It suffices to show that a.a.s. there are no connected bicyclic subgraph $M$ in $G_{D}[S]$ such that $W(M)$ has exactly $B$ vertices, $h$ vertices in $\overline{M}$ and $\ell$ vertices in $\underline{M}$. The number of ways to fix the set of vertices in $M$ and the vertices in $\overline{M}$ and $\underline{M}$ respectively is at most $O(n^{B})$. There are $O(1)$ ways to fix $W(M)$ on the set of chosen vertices. Finally, to form $M$, we consider all possible directed paths $P_{xy}$ where $(x,y)$ is an arc in $W(M)$. Hence, the probability that such a connected bicyclic subgraph $M$ exists in $G_{D}[S]$ is at most
@@
Proof of Claim 32 There are $a$ vertices, and every vertex has at most four neighbours. Specifying the set of neighbours for each vertex specifies a graph in ${\mathcal{F}}$. This immediately implies the upper bound $a^{4a}$. ∎
Proof of Claim 33 By definition of ${\mathcal{M}}$,
@@
Proof of Claim 34 Suppose that $w^{(i)}<4Kc_{n}$. Since $s_{y_{i}}\in S$, either down-set or the up-set of $s_{y_{i}}$ has size is at least $5Kc_{n}\alpha$ by (14). Since $|D_{P}[s_{y_{i}}]\cap M|\leq w^{(i)}+1$ and $|U_{P}[s_{y_{i}}]\cap M|=1$, it follows immediately that either $|D_{P}[s_{y_{i}}]\setminus M|$ has size at least $5Kc_{n}\alpha-4Kc_{n}\geq 4Kc_{n}\alpha$, or $|U_{P}[s_{y_{i}}]\setminus M|$ has size at least $5Kc_{n}\alpha-1\geq 4Kc_{n}\alpha$.
@@
Proof of (17).  It is obvious that $\mathcal{E}_{1}$ and $\mathcal{E}_{3}$ are independent and thus $\mathbb{P}(\mathcal{E}_{1}\cap\mathcal{E}_{3})=\mathbb{P}(\mathcal{E}_{1})\mathbb{P}(\mathcal{E}_{3})$. By Corollary 30, $\mathbb{P}(\mathcal{E}_{3})\leq c_{n}e^{c_{n}}n^{-1}$. The probability $\mathcal{E}_{1}$ is obviously maximized when $U$ is the largest $|U|$ elements in $[n]$ in the natural order. Moreover, since $|U|\leq 4Kc_{n}-1$ and $|D_{P}[U]|\geq(4Kc_{n}-1)\alpha$, at least one of the element in $U$ must have a down-set of size at least $\alpha$. Hence,
@@
by Corollary 30.
Proof of (18).  Let $P^{\prime}$ be the sub-poset induced by the elements in $[n]\setminus M$. Then,
$\mathbb{P}(\mathcal{E}_{2}\cap\mathcal{E}_{3})=\sum_{D}\mathbb{P}(D_{P^{\prime}}[s_{y_{i}}]=D)\mathbb{P}(\text{there is a directed arc joining $y_{i}$ and $D\cup\{s_{y_{i}}\}$}),$

### v2

@@ 6 Non-bipartite POS: proof of Theorem 12
Throughout this section, $p=c_{n}/n$ where $1\leq c_{n}\ll\log n$ and q=1pq=1-p. Set
@@
Proof of Theorem 12. By considering the dual of $P$ and Lemma 30, a.a.s. every component of the cover graph of $P[U[S]]$ is either a tree or unicyclic. By Theorems 3 and 4, a.a.s. $\dim P[D[S]],\dim P[U[S]]\leq 4$. Then, Theorem 8 implies that a.a.s. $\dim P\leq 2(8+\dim P[U\cup D]+\dim P[[n]\setminus S])$, where $U=U[S]\setminus S$ and $D=D[S]\setminus S$. By the definition of $S$, in both $P[U\cup D]$ and $P[[n]\setminus S]$, every element is comparable to at most $10Kc_{n}\alpha=O(c_{n}^{2}e^{c_{n}})$ other elements. Thus, by Theorem 5, $\dim P[U\cup D],\dim P[[n]\setminus S]=O(c_{n}^{4}e^{c_{n}})$. ∎
It only remains to prove Lemma 30. Unlike $P_{\mathcal{B}(n,n,p)}$ where the comparable graph and the cover graph are both $\mathcal{B}(n,n,p)$, the cover graph of $P_{\mathcal{G}(n,p)}$ is typically a proper subgraph of $\mathcal{G}(n,p)$, while the comparable graph of $P_{\mathcal{G}(n,p)}$ is a super graph of $\mathcal{G}(n,p)$ with a much greater density. To study $P_{\mathcal{G}(n,p)}$, it is much more convenient to consider the oriented version of $\mathcal{G}(n,p)$ where every edge $ij$ in $\mathcal{G}(n,p)$ is oriented from $i$ to $j$ if $i<j$. Since every graph on $[n]$ has a unique orientation we keep the same notation $\mathcal{G}(n,p)$ for its oriented version. Note that two elements $i$ and $j$ in $P_{\mathcal{G}(n,p)}$ are comparable if and only if there is a directed path joining $i$ and $j$ in $\mathcal{G}(n,p)$.
The key idea in the proof of upper-bounding $\dim P_{\mathcal{B}(n,n,p)}$ is that $S_{A}$ is much smaller than $A$, and thus, it is not likely to join vertices in $S_{A}$ by vertices in $B$ into a connected bicyclic subgraph. Situations in $P_{\mathcal{G}(n,p)}$ are quite different, as elements with large down-sets or up-sets form clusters. Suppose $v$ is an element with a large down-set. Then all the elements $u\in U_{P}(v)$ have a large down-set. Therefore, we expect many large cliques in the comparable graph of $P_{\mathcal{G}(n,p)}$, and they intersect in non-trivial ways. However, this does not mean that the cover graph of $P_{\mathcal{G}(n,p)}$ induced by the elements with large down-sets have complicated graph structures. Indeed, Lemma 30 shows that every component of the cover graph has at most one cycle. Unlike the study of $\dim P_{\mathcal{B}(n,n,p)}$, analysing the cover graph of $P[D[S]]$ for $P\sim P_{\mathcal{G}(n,p)}$ necessitates the identification and careful examination of suitable chosen structures.
@@
Proof of Observation 31. Suppose on the contrary that there exists an arc $(x,y)$ in $G_{D}[S]\setminus\mathcal{G}(V,p)$. It implies that there is a directed $x,y$-path $P_{xy}$ in $\mathcal{G}(n,p)$ such that $P_{xy}$ has length at least two, and none of the internal vertices is in $V$. On the other hand, there is a directed path $P_{yz}$ from $y$ to some $z\in S$ since $y$ is in the down-set of $S$. By considering the directed path obtained from concatenating $P_{xy}$ and $P_{yz}$, it follows that all the internal vertices on $P_{xy}$ must be in the down-set of $z$, and thus must be in $V$, leading to a contradiciton. ∎
It is easy to see that the cover graph of $P[D[S]\cup U[S]]$ does not have the property in Observation 31, which is one of the main reasons that we complete the proof of Theorem 12 by studying the dimension of $P[D[S]]$ and $P[U[S]]$, instead of studying $\dim P[D[S]\cup U[S]]$.
@@
To simplify the analysis, we hope to identify a set of directed paths as in Observation 32(b) that “behave somewhat independently”. We say the triple $(x,y,z)$ in $W(M)$ is a trilink if
@@
Choose $M$ such that the total number of maximal elements is minimized. Let $\mathcal{T}$} be a maximum collection of pairwise well-separated trilinks in $W(M)$, and for each trilink $(x,y,z)\in{\mathcal{T}}$, let $P_{y}$ be an arbitrary directed path from $y$ to some $s_{y}\in S$, whose existence is guaranteed by Observation 32(b). Let $\mathcal{P}$} be the collection of such directed paths. We demonstrate that our choice of $(M,{\mathcal{T}},{\mathcal{P}})$ satisfies (P1)–(P4).
(P1): it holds trivially when $t\leq 42$ by taking $\mathcal{T}$$\emptyset$. Suppose that $t\geq 43$. Let $M^{\prime}$ be the vertex-induced subgraph of $W(M)$ obtained by deleting the red vertices in $W(M)$ and their neighbours. Since there were either two red vertices each of degree three, or a single red vertex with degree 4 in $W(M)$, it is easy to see that $M^{\prime}$ has at least t12t-12 edges. Moreover, $M^{\prime}$ is a collection of at most three disjoint paths (not necessarily directed paths), one of which — call it $P$ — must have length at least (t12)/3(t-12)/3. Since every edge in $M^{\prime}$ joins a vertex in $\overline{M}$ and a vertex in $\underline{M}$, there must exists at least $\lfloor((t-12)/3)/7\rfloor\geq t/21-2$ well-separated trilinks by taking a maximum collection of pairwise well-separated trilinks on $P$.
@@
Let $U\subseteq[n]$ and let $G_{\bar{U}}$ be the subgraph of $\mathcal{G}(n,p)$ induced by $n]\setminus U$. Let $P_{\bar{U}}$ be the poset with cover graph $G_{\bar{U}}$.
@@
Proof of Lemma 34. For (a), obviously, $\mathbb{P}(|D_{P_{\bar{U}}}(u)|\geq k)\leq\mathbb{P}(|D_{P}(u)|\geq k)$, and we can obtain the bound in part (a) easily from (6). However, it follows even more straightforwardly from [6, Lemma 3.1 (iii)].
@@
    For every $u\in[n]$, $\mathbb{P}(|D_{P}[u]\cup U_{P}[u]|\geq\alpha)\leq(1+o(1))\exp(-Kc_{n})$.
@@
Proof of Lemma 30. Suppose $G_{D}[S]$ has a connected bicyclic subgraph $M$ and a collection $\mathcal{P}$} of dipaths satisfying all the properties in Lemma 33. By the claim below, we may assume that $W(M)$ has at least 100 vertices.
@@
We first bound the number of ways to construct $W(M)$, and then $M$. We define a family $\mathcal{F}$} of connected directed bicyclic graphs (who can possibly contain a double edge) that can be candidates for $W(M)$ as follows. The vertices are marked ++, or -, or coloured red, so that they respond to the vertices in $\overline{M}$, $\underline{M}$, and the red vertices in $W(M)$.
@@
  • Each vertex can be marked ++ or -, or unmarked;
@@
  • Every arc $(x,y)$ joins some $x$ marked - and some $y$ marked ++, unless $x$ or $y$ is red.
@@
Let $\Sigma_{1},\Sigma_{2}$ be disjoint sets. Let $R$ be a set (not necessarily disjoint from $\Sigma_{1}\cup\Sigma_{2}$) of cardinality between one and two, and let $a=|\Sigma_{1}\cup\Sigma_{2}\cup R|$. The number of graphs in $\mathcal{F}$} on $\Sigma_{1}\cup\Sigma_{2}\cup R$ such that all vertices in $\Sigma_{1}$ are marked ++, all vertices in $\Sigma_{2}$ are marked -, and all vertices in $R$ are coloured red is at most $O(a^{4a})$.
@@
Explanation of (15) and the notation used therein. The first sum is over all integers $h,\ell,a$ that denote the sizes of $\Sigma_{1}$, $\Sigma_{2}$ and the total number of vertices in $W(M)$. The second sum is over all possible choices of $(\Sigma_{1},\Sigma_{2},R)$ that is consistent with $(h,\ell,a)$, where $R$ denotes the set of red vertices in $W(M)$. The third sum is over all possible weights on the arcs of $W(M)$, which indicates the lengths of the directed paths in $M$ between two vertices of $W(M)$. The next sum is over all possible choices of $M$ that are consistent with $(\Sigma_{1},\Sigma_{2},R,{\boldsymbol{w}})$. The next sum is over all choices of $k$ well-separated trilinks. Given a set of trilink{{trilink}}s $\mathcal{T}$}, let ${\mathcal{N}}={\mathcal{N}}({\boldsymbol{w}})$ be the set of $j\in[k]$ such that $w^{(j)}<4Kc_{n}$ where $w^{(j)}$ is for $y_{j}$ where $(x_{i},y_{i},z_{i})\in{\mathcal{T}}$. Next, we sum over all choices of $s_{y_{j}}$ for $j\in{\mathcal{N}}({\boldsymbol{w}})$. The final sum is over all possible subsets $D_{j},U_{j},j\in{\mathcal{N}}$ that are pairwise disjoint, except that $D_{j}\cap U_{j}=y_{j}$ for all $j$. Recalling that $N({\boldsymbol{w}})=|{\mathcal{N}}({\boldsymbol{w}})|=|\{i:w^{(i)}<4Kc_{n}\}|$. By Claim 39(a), for each $j\in{\mathcal{N}}({\boldsymbol{w}})$ either $|D_{j}|\geq 4Kc_{n}\alpha$ or $|U_{j}|\geq 4Kc_{n}\alpha$. Thus, by Claim 39(b,c),
@@
$\hskip 227.62204pt\times n^{N}2^{a}\Big(2n^{-1}\exp(-Kc_{n}/2)\Big)^{N}.$ (16)
Explanation of (16). We call $j\in[a+1]$ heavy if $w_{j}\geq 2Kc_{n}$ and light otherwise. Since $N({\boldsymbol{w}})=N$, there are kNk-N $y_{i}$s in a trilink in $\mathcal{T}$} such that $w^{(i)}\geq 4Kc_{n}$. Since each of these $y_{i}$s has degree two in $W(M)$, at least one of the directed path incident to $y_{i}$ in $M$ has length at least $2Kc_{n}$, i.e. the corresponding $j$ is heavy. Since each directed path joins two vertices in $M$, there must be at least (kN)/2(k-N)/2 heavy $j\in[a+1]$. Let $\mathcal{A}$} be the set of all possible assignments of “heavy” and “light” to $j\in[a+1]$ such that there are at least (kN)/2(k-N)/2 heavy elements. Obviously, $|{\mathcal{A}}|\leq 2^{a+1}$. Then,
@@
by choosing sufficiently large $K$ so that $8c_{n}e^{c_{n}-Kc_{n}/168}<1/2$. ∎
Proof of Claim 36. Fix $B,h,\ell$. It suffices to show that a.a.s. there are no connected bicyclic subgraph $M$ in $G_{D}[S]$ such that $W(M)$ has exactly $B$ vertices, $h$ vertices in $\overline{M}$ and $\ell$ vertices in $\underline{M}$. The number of ways to fix the set of vertices in $M$ and the vertices in $\overline{M}$ and $\underline{M}$ respectively is at most $O(n^{B})$. There are $O(1)$ ways to fix $W(M)$ on the set of chosen vertices. Finally, to form $M$, we consider all possible directed paths $P_{xy}$ where $(x,y)$ is an arc in $W(M)$. Hence, the probability that such a connected bicyclic subgraph $M$ exists in $G_{D}[S]$ is at most
@@
Proof of Claim 37. There are $a$ vertices, and every vertex has at most four neighbours. Specifying the set of neighbours for each vertex specifies a graph in $\mathcal{F}$}. This immediately implies the upper bound $a^{4a}$. ∎
Proof of Claim 38. By definition of $\mathcal{M}$},
@@
Proof of Claim 39. Suppose that $w^{(i)}<4Kc_{n}$. Since $s_{y_{i}}\in S$, either down-set or the up-set of $s_{y_{i}}$ has size is at least $5Kc_{n}\alpha$ by (14). Since $|D_{P}[s_{y_{i}}]\cap M|\leq w^{(i)}+1$ and $|U_{P}[s_{y_{i}}]\cap M|=1$, it follows immediately that either $|D_{P}[s_{y_{i}}]\setminus M|$ has size at least $5Kc_{n}\alpha-4Kc_{n}\geq 4Kc_{n}\alpha$, or $|U_{P}[s_{y_{i}}]\setminus M|$ has size at least $5Kc_{n}\alpha-1\geq 4Kc_{n}\alpha$.
@@
Proof of (17). It is obvious that $\mathcal{E}_{1}$ and $\mathcal{E}_{3}$ are independent and thus $\mathbb{P}(\mathcal{E}_{1}\cap\mathcal{E}_{3})=\mathbb{P}(\mathcal{E}_{1})\mathbb{P}(\mathcal{E}_{3})$. By Corollary 35 (b), $\mathbb{P}(\mathcal{E}_{3})\leq c_{n}e^{c_{n}}n^{-1}$. The probability $\mathcal{E}_{1}$ is obviously maximized when $U$ is the largest $|U|$ elements in $[n]$ in the natural order. Moreover, since $|U|\leq 4Kc_{n}-1$ and $|D_{P}[U]|\geq(4Kc_{n}-1)\alpha$, at least one of the element in $U$ must have a down-set of size at least $\alpha$. Hence,
@@
by Corollary 35 (a).
Proof of (18). Let $P^{\prime}$ be the sub-poset induced by the elements in $n]\setminus M$. Then,
[…]